# Single-Channel EEG Tokenization Through Time-Frequency Modeling

### v1

@@ 3.1. Notations and Problem Statements
Remark. We here hold several expectations for the learned motif tokens. First, these tokens are expected to reduce redundancy, noise, and complexity, providing a compact, sparse, and informative representation of EEGs. Second, these motifs should effectively capture essential neurophysiological patterns from both temporal and frequency domains. Third, the tokens should generalize well across different EEG tasks, enhancing the efficiency and interpretability of the data. We set up related research questions to evaluate these expectations in Section. 5.
Problem Statement 2 (Multi-Channel EEG Classification): Given EEGs $\mathbf{X}$ and a fixed, learned single-channel tokenizer $f_{\text{tokenizer}}$, we apply $f_{\text{tokenizer}}$ independently to each channel $c$ to obtain a tokenization representation $\Bigl{\{}\{v_{i}^{c}\}_{i=1}^{N}\Bigr{\}}_{c=1}^{C}$. Then, these tokens can serve various downstream tasks. For classification tasks, they are mapped to labels by:
$f_{\text{classifier}}:(\mathcal{V}^{D})^{N\times C}\rightarrow\mathbf{Y}$
where $Y$ is the target labels (e.g., EEG events, seizure types). By aggregating and processing the tokens across all channels, $f_{\text{classifier}}$ predicts a label $y\in\mathbf{Y}$. Notably, $f_{\text{classifier}}$ can be any downstream-oriented model, and its training is performed separately from the EEG tokenizer $f_{\text{tokenizer}}$.
4. Methodology
TFM-Token comprises two components: (1) TFM-Tokenizer: addresses problem statement 1 by converting continuous EEG signals into discrete tokens, capturing key time-frequency features, and (2) TFM-Encoder: tackles problem statement 2 by leveraging these tokens for downstream EEG tasks, such as classification. To mitigate the quadratic complexity of standard Transformers (Vaswani, 2017), we employ a linear attention mechanism (Katharopoulos et al., 2020; Wang et al., 2020) across all transformer modules. Given input EEGs, TFM-Tokenizer independently tokenizes each channel into discrete token sequences. The TFM-Encoder then processes these tokens to perform classification. For training TFM-Token, we first conduct an unsupervised pretraining of TFM-Tokenizer in a single-channel setting to encode time-frequency motifs into discrete tokens (see Figure 2a, Sec 4.1). The tokenizer is then frozen, and TFM-Encoder undergoes masked token prediction pretraining (see Figure 2b, Sec 4.2), followed by fine-tuning for downstream EEG tasks.
Time-Frequency discretization. The input to TFM-Token consists of discrete time and frequency patches, applied to both raw EEG data and its corresponding spectrogram. This approach addresses the transient nature of EEG waveforms, which may decompose the entangled patterns and is foundational for motif learning. Specifically, we segment $x$ into time-domain length patches $L$ with the same hop size of $H$. This yields a sequence of patches $\{x_{i}\}_{i=1}^{N}$, where $N=\lfloor(T-L)/H\rfloor+1$ is the total number of patches. Using the same $L$ and $H$ for both STFT and temporal segmentation, we ensure that each patch $x_{i}$ corresponds one-to-one with a spectral window $\mathbf{S}_{i}$. In practice, we obtain the raw EEG patches by configuring the convolutional layers with a kernel size and stride that match the STFT window length $L$ and hop size $H$. Additionally, we denote embeddings and masks as $\mathbf{E}$ and $\mathbf{M}$, respectively.
4.1. Single Channel TFM-Tokenizer
We introduce the TFM-Tokenizer, a scalable module for tokenizing single-channel EEG signals $x$ by effectively capturing both their temporal and frequency dynamics. Our design is inspired by the Vector-Quantized Variational Autoencoder (VQ-VAE)(Van Den Oord et al., 2017), which has been widely adopted for tokenization efforts in other domains such as video processing (Agarwal et al., 2025). As illustrated in Figure 2a, the TFM-Tokenizer comprises three primary components: (1) Localized Spectral Window Encoder, (2) Temporal Encoder, and (3) Temporal Transformer. At a high level, TFM-Tokenizer adopts a frequency-then-time paradigm. First, we isolate each spectral window $\mathbf{S}_{i}$ (for $i=1,…,N$) to capture intra-frequency band dependencies, yielding an aggregated representation for each window (Localized Spectral Window Encoder, see Figure 2d). Next, these frequency-based representations are combined with features extracted from the corresponding raw EEG patches $\{x_{i}\}_{i=1}^{N}$ (Temporal Encoder). Finally, the combined features are passed through the Temporal Transformer to capture long-range temporal dependencies, and the resulting embeddings are quantized into discrete tokens.
Localized Spectral Window Encoder. Capturing frequency-band characteristics or compositions is crucial for EEG analysis, as the data often contains oscillatory components (e.g., alpha, beta) with varying amplitudes and time. Unlike previous studies that project an entire spectral window with a single linear layer (Yang et al., 2024), we propose to patch along the frequency axis, enabling better modeling of cross-frequency dependencies. In detail, this encoder consists of three main steps (see Figure 2d):
  • Frequency Patch Encoder. Given a set of spectral windows $\{\mathbf{S}_{i}\}_{i=1}^{N}$, we isolate and divide each spectral window $\mathbf{S}_{i}$ into $P$ non-overlapping patches $\{\mathbf{S}_{(i,p)}\}_{p=1}^{P}$, each spanning $\Delta f$ frequency bins such that $P.\Delta f=F$. We then project each frequency patch into a latent space:
    $e_{(i,p)}=\text{GroupNorm}\left(\text{GeLU}\left(\mathbf{W}_{p}\mathbf{S}_{(i,p)}\right)\right)$
    where $\mathbf{W}_{p}\in\mathbb{R}^{D\times\Delta f}$ is a learnable matrix.
  • Frequency Transformer. In order to capture the interactions among the frequency patch embeddings $e_{(i,p)}$, we apply a frequency transformer that operates along the frequency axis within each spectral window $\mathbf{S}_{i}$. This enables the capturing of intra-frequency band dependencies (e.g., interactions between low and high-frequency bands).
  • Gated Patchwise Aggregation: In many EEG scenarios, large portions of the frequency spectrum can be irrelevant. For instance, tasks related to sleep primarily focus on frequency bands up to approximately 32 Hz (Chen et al., 2023). Also, the frequencies of interest vary across conditions and tasks. To emphasize important frequency patches and suppress the rest, we adopt a gated aggregation mechanism. Formally, we first group frequency patch embeddings, then apply gated weighting before concatenating them into a single embedding $\mathbf{E}^{F}_{i}$
    $\mathbf{E}^{F}_{i}=\text{Concat}\left[\sigma\left(\mathbf{W_{g1}e_{(i,p)}}\right)\mathbf{W_{g2}e_{(i,p)}}\right]$
    where $\mathbf{W_{g1}},\mathbf{W_{g2}}$ are trainable parameters and $\sigma(\cdot)$ is the element-wise sigmoid function.
Temporal Encoder. To capture temporal dynamics from the raw EEG patches $\{x_{i}\}_{i=1}^{N}$, we perform a linear projection of the patch followed by GELU(Hendrycks and Gimpel, 2016) activation and group normalization, producing temporal embeddings $\{\mathbf{E}^{T}_{i}\}_{i=1}^{N}$.
Temporal Transformer. We form a unified time-frequency embedding by combining each aggregated frequency embedding $\mathbf{E}_{i}^{F}$ with its corresponding temporal embedding $\mathbf{E}_{i}^{T}$ and feed the resulting vectors into a temporal transformer. This module attends jointly to time and frequency features across all windows, capturing long-range dependencies. The output is then quantized into discrete tokens $\{v_{i}\}_{i=1}^{N}$ using a learnable codebook $\mathcal{V}^{k}$. Notably, we do not add any positional encoding as EEG signals are inherently non-stationary and often exhibit chaotic behavior; thus, our goal is to capture their distinctive features without enforcing positional constraints. We further provide an ablation study (Appendix A.3) showing the effect of omitting positional encodings.
Tokenizer Codebook. Our tokenizer is designed to capture temporal frequency motifs by applying vector quantization along the time axis, that is, individual patches. This is different from conventional quantization methods in computer vision, which typically operate on the embedding dimension of a patch (Van Den Oord et al., 2017; Esser et al., 2020). As a result, each token in our vocabulary directly represents a waveform within a specific short duration and can facilitate the retrieval of the timestamp of input EEGs to understand the composition of an EEG. This enhances interpretability; for instance, specific codes in the vocabulary are activated in response to the occurrence of particular temporal or frequency patterns. To validate this, we conducted a visual inspection study, with results presented in Section 5.5.
Frequency Making Prediction for Tokenizer Learning. The input to the tokenizer is the full frequency band of a temporal patch. To facilitate effective frequency learning, we employ a frequency-band and temporal masking strategy during TFM-Tokenizer training. We partition $\mathbf{S}$ into $N_{F}=\lfloor\frac{F}{\delta_{f}}\rfloor$ frequency groups of size $\delta_{f}$ and randomly mask a subset based on a predefined mask ratio (e.g., 0.5), forming a frequency-band mask $M_{F}$. Simultaneously, a temporal mask $M_{T}$ is applied along the time axis. The final mask, $M=M_{F}\land M_{T}$, is used to mask the STFT representation, $S^{M}=M\odot S$, where $\odot$ denotes element-wise multiplication. For data augmentation and efficient training, we adopt a symmetric masking approach from (Jiang et al., 2024) by defining $M_{\text{sym}}=1-M$. Given $\mathbf{S}^{M}$ and $x$, our TFM-Tokenizer first quantized them into a sequence of distinct tokens, which is then passed through a transformer decoder followed by a linear decoder that predicts the masked regions. The tokenizer is trained using a masked reconstruction loss:
$\mathcal{L}_{\mathrm{rec}}=\sum_{(f,t)\in\mathcal{I}_{\mathrm{mask}}}\!\bigl{\|}\mathbf{S}(f,t)-\hat{\mathbf{S}}(f,t)\bigr{\|}_{2}^{2}$
where $(f,t)\in\mathcal{I}_{\text{mask}}$ represents masked frequency–time indices and $\hat{\mathbf{S}}$ is the reconstruction of $\mathbf{S}$. Alongside the masked reconstruction loss, we incorporate the codebook and commitment losses (Van Den Oord et al., 2017), yielding the total loss:
$\mathcal{L}_{\mathrm{token}}=\mathcal{L}_{\mathrm{recon}}+\alpha\;\sum_{i}\bigl{\|}\mathrm{sg}[E_{i}]\;-\;v_{i}\bigr{\|}_{2}^{2}+\beta\;\sum_{i}\bigl{\|}E_{i}\;-\;\mathrm{sg}[v_{i}]\bigr{\|}_{2}^{2}$
where, $\mathrm{sg}[\cdot]$ is the stop-gradient operator, and $\alpha,\beta$ are hyperparameters. We also employ exponential moving average strategy(Van Den Oord et al., 2017) for stable codebook updates.
4.2. Token-Wise TFM-Encoder
The TFM-Encoder is designed to aggregate tokenized representations across channels and perform downstream classification. It consists of two main components: (1) a token-embedding lookup table and (2) linear attention transformer layers. Given a multi-channel recording $\mathbf{X}\in\mathbb{R}^{C\times T}$, we apply the fixed (pretrained) TFM-Tokenizer to each channel independently, obtaining discrete token sequences $\Bigl{\{}\{v_{i}^{c}\}_{i=1}^{N}\Bigr{\}}_{c=1}^{C}$. Each token is mapped to its embedding via a lookup table initialized with the TFM-Tokenizer ’s learned codebook weights. We flatten the embeddings across channels and incorporate channel and position embeddings to encode respective information. We then prepend a [CLS] token (Devlin, 2018), the output of which is designed to capture the overall representation of the sequence for downstream tasks. The resulting embeddings are then processed by linear attention transformers in TFM-Encoder.
Token Prediction for TFM-Encoder Learning. We adopt a strategy akin to masked language modeling. We randomly mask tokens across multiple channels and time steps and train the model to predict these masked tokens via a cross-entropy loss. Along with representation learning, this approach enhances robustness to missing or corrupted data, common in real-world EEG systems where channels or time segments may be dropped or noisy. Finally, the TFM-Encoder is finetuned to downstream tasks.
5. Experiments and Results
Table 1. EEG classification performance comparison on TUEV and TUAB datasets.
Models Number TUEV (event type classification) TUAB (abnormal detection)
of Params Balanced Acc. Cohen’s Kappa Weighted F1 Balanced Acc. AUC-PR AUROC
SPaRCNet(Jing et al., 2023) 0.79M $0.4161\pm 0.0262$ $0.4233\pm 0.0181$ $0.7024\pm 0.0104$ $0.7885\pm 0.0205$ $0.8782\pm 0.0127$ $0.8804\pm 0.0071$
ContraWR(Yang et al., 2023) 1.6M $0.4384\pm 0.0349$ $0.3912\pm 0.0237$ $0.6893\pm 0.0136$ $0.7851\pm 0.0050$ $0.8719\pm 0.0076$ $0.8743\pm 0.0042$
CNN-Transformer(Peh et al., 2022) 3.2M $0.4087\pm 0.0161$ $0.3815\pm 0.0134$ $0.6854\pm 0.0293$ $0.7898\pm 0.0062$ $0.8781\pm 0.0053$ $0.8781\pm 0.0067$
FFCL(Li et al., 2022) 2.4M $0.3979\pm 0.0104$ $0.3732\pm 0.0188$ $0.6783\pm 0.0120$ $0.7869\pm 0.0054$ $0.8694\pm 0.0091$ $0.8761\pm 0.0044$
ST-Transformer(Song et al., 2021) 3.5M $0.3984\pm 0.0228$ $0.3765\pm 0.0306$ $0.6823\pm 0.0190$ $0.8004\pm 0.0037$ $0.8737\pm 0.0069$ $0.8798\pm 0.0029$
BIOT(Yang et al., 2024) 3.2M $0.4679\pm 0.0354$ $0.4890\pm 0.0407$ $0.7352\pm 0.0236$ $0.7955\pm 0.0047$ $0.8819\pm 0.0046$ $0.8834\pm 0.0041$
LaBraM-Base(Jiang et al., 2024) 5.8M $0.4682\pm 0.0856$ $0.5067\pm 0.0413$ $0.7466\pm 0.0202$ $0.7720\pm 0.0046$ $0.8498\pm 0.0036$ $0.8534\pm 0.0027$
TFM-Token-R 1.8M $\underline{0.4898}\pm 0.0105$ $0.5194\pm 0.0195$ $0.7518\pm 0.0095$ $\underline{0.8033}\pm 0.0021$ $\underline{0.8908}\pm 0.0027$ $\underline{0.8849}\pm 0.0024$
TFM-Token-S 1.9M $0.4708\pm 0.0339$ $\underline{0.5275}\pm 0.0314$ $\underline{0.7538}\pm 0.0152$ $0.7927\pm 0.0044$ $0.8814\pm 0.0095$ $0.8836\pm 0.0052$
TFM-Token 1.9M 0.4943 $\pm$ 0.0516 0.5337 $\pm$ 0.0306 0.7570 $\pm$ 0.0163 0.8152 $\pm$ 0.0014 0.8946 $\pm$ 0.0008 0.8897 $\pm$ 0.0008
1. The best results are bolded, while the second-best are underlined. 2. The number of parameters for LaBraM is only considering their classifier model. The size of their neural tokenizer was 8.6M. 3. Results on CHB-MIT and IIIC Seizure are provided in Table 4 and 5 in Appendix A.1
In this section, we perform comprehensive experiments to evaluate our proposed method to answer the following research questions:
  • Q1: Can a fully discretized EEG tokenization framework outperform baseline methods that rely on continuous embeddings? How crucial is joint frequency-temporal modeling in EEG analysis and tokenization?
  • Q2: How effectively do our learned tokens represent EEG features and capture class-specific characteristics? Do the generated tokens capture relevant frequency information?
  • Q3: Are the tokens generated by our TFM-Tokenizer scalable, and can they enhance the performance of existing EEG foundation models such as LaBraM?
  • Q4: Do tokens produced by TFM-Tokenizer offer interpretability, and do they correspond to distinct, recognizable EEG patterns?
5.1. Experiment Setup
5.1.1. Datasets:
We conducted our evaluation on four EEG datasets:
  • TUH EEG Events (TUEV) (Harati et al., 2015): TUEV is a subset of the TUH EEG Corpus (Obeid and Picone, 2016), which comprises clinical EEG recordings collected at Temple University Hospital between 2002 and 2017. The dataset is annotated for six EEG event types: spike and sharp wave (SPSW), generalized periodic epileptiform discharges (GPED), periodic lateralized epileptiform discharges (PLED), eye movement (EYEM), artifact (ARTF), and background (BCKG).
  • TUH Abnormal EEG Corpus (TUAB) (Lopez et al., 2015): TUAB comprises EEG recordings collected at Temple University Hospital, which are labeled for normal and abnormal EEG activity.
  • IIIC Seizure (Jing et al., 2023; Ge et al., 2021): The IIIC Seizure dataset is curated for the detection of six distinct ictal–interictal–injury continuum (IIIC) patterns and is sourced from (Jing et al., 2023; Ge et al., 2021). The annotations include: (1) others (OTH), (2) seizure types (ESZ), (3) lateralized periodic discharge (LPD), (4) generalized periodic discharge (GPD), (5) lateralized rhythmic delta activity (LRDA), and (6) generalized rhythmic delta activity (GRDA).
  • CHB-MIT (Shoeb, 2009): The CHB-MIT dataset is a widely used benchmark for epilepsy seizure detection. It comprises EEG recordings from 23 pediatric subjects with intractable seizures.
5.1.2. Preprocessing:
@@
We follow the preprocessing setup of BIOT (Yang et al., 2024). Unlike LaBraM (Jiang et al., 2024), which utilized 23 channels in the TUEV and TUAB datasets, we adhere to the 16-channel bipolar montage from the international 10–20 system, as used in (Yang et al., 2024). All EEG recordings are resampled to 200 Hz. For TUEV and TUAB, we apply a bandpass filter (0.10.10.10.175757575 Hz) and a notch filter (50 Hz), following the preprocessing pipeline of LaBraM (Jiang et al., 2024). STFT computation of the signals is performed using PyTorch, with detailed parameters provided in Appendix C.2. For training, validation, and test splits, we follow the recommendations from (Yang et al., 2024). Additional details on dataset statistics and splits are provided in Appendix C.1.
5.1.3. Baselines and Metrics:
We evaluated our approach against the baselines from (Yang et al., 2024) as well as the current state-of-the-art methods, including BIOT (Yang et al., 2024) and LaBraM (Jiang et al., 2024). All baselines were reproduced using their respective open-source GitHub repositories. To ensure a fair comparison, our experiments follow a single-dataset setting for all the baselines. Specifically for BIOT, we conducted their proposed unsupervised pretraining followed by fine-tuning on the same dataset. Similarly, for LaBraM, we used their base model and conducted neural tokenizer training, masked EEG modeling, and fine-tuning within the same dataset. For performance evaluation, we used balanced accuracy, Cohen’s Kappa coefficient, and weighted-F1 score for multi-class classification tasks, while balanced accuracy, AUC-PR, and AUROC were used for binary classification tasks. For TUAB, we used binary cross-entropy loss for fine-tuning, while the cross-entropy loss was applied to the TUEV and IIIC datasets. Given the class imbalance in the CHB-MIT dataset, we employed focal loss for all experiments. All experiments were conducted using five different random seeds, and we report the mean and standard deviation for each metric. Appendix C provides additional details on the experiment settings.
5.2. Q1: Performance and Importance of Joint Frequency-Temporal Modeling
Q1.1 - Performance Evaluation: Table 1 presents EEG event classification results on TUEV and abnormal detection performance on TUAB. Our TFM-Token consistently outperforms all baselines across all metrics. For event-type classification on TUEV, TFM-Token achieves a $5\%$ increase in balanced accuracy over BIOT ($0.4679\rightarrow 0.4943$) and LaBraM ($0.4682\rightarrow 0.4943$). In Cohen’s Kappa, TFM-Token improves by $9\%$ over BIOT ($0.4890\rightarrow 0.5337$) and $5\%$ over LaBraM ($0.5067\rightarrow 0.5337$). In abnormal detection on TUAB, TFM-Token achieves a $5\%$ improvement across all metrics compared to LaBraM. This highlights that our fully discrete tokenization approach surpasses the performance of existing continuous embedding-based approaches. Another advantage of TFM-Token is its reduced model footprint. As shown in Table 1, TFM-Token achieves better results with significantly fewer parameters—a 3-fold reduction compared to LaBraM (5.8M $\rightarrow$ 1.9M) and a 1.5-fold reduction compared to BIOT (3.2M $\rightarrow$ 1.9M). This reduction can be attributed to discrete tokenization approach, which compresses EEG into a token sequence, thereby reducing data complexity.
Q1.2 - Importance of Joint Frequency and Temporal Modeling: To evaluate the importance of joint frequency-temporal modeling, we conducted an ablation study comparing three tokenization variants: (1) TFM-Token-Raw Signal Only (TFM-Token-R), which uses only raw EEG patches $\{x_{i}\}_{i=1}^{N}$ to predict the spectrum $\mathbf{S}$, (2) TFM-Token-STFT Only (TFM-Token-S), and (3) TFM-Token, which jointly models both temporal and frequency features. Masked modeling was applied for token learning in the latter two, with consistent TFM-Encoder training across all variants. As shown in Table 1, all three variants outperform existing baselines. In event classification, TFM-Token-S improves Cohen’s Kappa over TFM-Token-R ($0.5194\rightarrow 0.5275$). However, in abnormal detection, TFM-Token-R achieves a higher AUC-PR ($0.8814\rightarrow 0.8908$). These results indicate that different EEG tasks rely on distinct feature domains, underscoring the necessity of joint modeling. The primary TFM-Token consistently outperforms both single-domain approaches across all settings, further underscoring the importance of joint modeling.
5.3. Q2: EEG Token Quality Analysis and Frequency Learning
Table 2. Token Utilization and class-token uniqueness comparison

### v2

@@ Appendix A Problem Formulation
Problem Statement 1 (EEG Tokenization): Given a single channel EEG $x$, we aim to learn a tokenization function $f_{\textbf{tokenizer}}:\mathbb{R}^{T}\rightarrow\mathcal{V}^{N\times D}$, that maps $x$ (or transformations) to a sequence of discrete tokens $\{v_{i}\}_{i=1}^{N}$, where each token is from a learnable EEG token vocabulary $\mathcal{V}$ of size $k$ and embedding size of $D$. These tokens should represent various time-frequency “motifs” derived from both $x_{i}$} and $\mathbf{S}_{i}$}. Therefore, $\mathcal{V}$ is learnable from $\mathbf{S}$ and the temporal patches $\{x_{i}\}_{i=1}^{N}$. Remark. We here hold several expectations for the learned motif tokens. First, these tokens are expected to reduce redundancy, noise, and complexity, providing a compact, sparse, and informative representation of EEGs. Second, these motifs should capture key neurophysiological patterns from both temporal and frequency domains. Third, the tokens should generalize well across different EEG tasks.
[…]
Appendix B Additional Experiment Details
B.1 Dataset Statistics and Splits
Table 4: Evaluation Dataset Summary
Dataset # of Recordings # of Samples Duration (s) Task
TUEV $11,914$ $112,491$ 55 EEG Event Classification
IIIC Seizure $2,689$ $135,096$ $10$ Seizure Type Classification
CHB-MIT $686$ $326,993$ $10$ Seizure Detection
TUAB $2,339$ $409,455$ $10$ Abnoral EEG Detection
EESM23 $120$ $14,509$ $30$ Ear-EEG based Sleep Staging
This section provides detailed information on the datasets used in our experiments and their respective splits. Table 4 summarizes key statistics, including the number of recordings, the total number of samples after preprocessing, their duration, and the corresponding downstream tasks. For TUEV and TUAB, we utilized the official training and test splits provided by the dataset and further divided the training splits into $80\%$ training and $20\%$ validation sets. We performed a subject-wise split into $60\%$ training, $20\%$ validation, and $20\%$ test on the IIIC Seizure dataset. In the CHB-MIT dataset, we used 1-19 subjects for training, 20-21 for validation, and 22-23 for testing. For the out-of-distribution evaluation on the ear-EEG EESM23 (Bjarke Mikkelsen et al., 2025) dataset, we followed a subject-wise split, where subjects 1–6 were used for fine-tuning, 7–8 for validation, and 9–10 for testing.
B.2 Preprocessing
@@
We follow the preprocessing setup of BIOT  (Yang et al., 2024). We adhere to the 16-channel bipolar montage from the international 10–20 system, as used in  (Yang et al., 2024). All EEG recordings are resampled to 200 Hz. For TUEV and TUAB, we apply a bandpass filter (0.10.17575 Hz) and a notch filter (50 Hz), following the preprocessing pipeline of LaBraM  (Jiang et al., 2024b). We then segment the recordings according to the provided annotations and preprocessing guidelines. STFT computation of the signals is performed using PyTorch, with detailed parameters provided in Appendix B.6. For training, validation, and test splits, we follow the recommendations from  (Yang et al., 2024). We adopt a window length of 11s with $0.5$s overlap to segment EEG signals during training and inference, following prior work for consistency (Yang et al., 2024).
B.3 Ear-EEG Preprocessing
We follow the preprocessing guidelines of Tabar et al. (2021) for the EESM-23 ear-EEG dataset, which includes four channels (RB, RT, LB, LT). A bandpass filter (0.10.1100100 Hz) and a 50Hz notch filter are applied. Each patients perform certain tasks before sleep. To isolate sleep segments, we crop each session from the onset of annotated sleep scoring, segment the signal into 30-second epochs, and discard corrupted segments.
B.4 Evaluation Metrics
For evaluation, we used balanced accuracy, Cohen’s kappa coefficient, and weighted F1 for multi-class classification, and balanced accuracy, AUROC, and AUC-PR for binary classification. During finetuning, we employed binary cross-entropy loss for TUAB, cross-entropy loss for TUEV and IIIC, and focal loss for CHB-MIT due to class imbalance. All experiments were conducted using five different random seeds, and we report the mean and standard deviation.
B.5 Additional details on baselines
All baselines were reproduced using their official open-source repositories. LaBraM’s primary contribution lies in large-scale EEG pretraining using over 2,500 hours of data (Jiang et al., 2024b), whereas our focus is on developing an effective EEG tokenizer. To ensure a fair comparison, we reproduced LaBraM using its official repository under our dataset and experimental settings. For EEGPT, we report the published results for the 4.7M model on TUEV and TUAB (Wang et al., 2024a). Since results on CHB-MIT and IIIC-Seizure were not available, we used the official pretrained weights and fine-tuned the model on these tasks.
B.6 STFT parameters
Table 5: STFT parameters
Parameter Value Description
Hop length $H$ $100$ Step size for sliding window ($50\%$ overlap)
Centering False The STFT is computed without implicit zero-padding
To extract frequency-domain representations of the EEG, we utilized the STFT function from PyTorch. The recommendations of Yang et al. (2024) guided our parameter selection and empirical analysis of different configurations to optimize the trade-off between time-frequency resolution. The final parameters are as follows:
Appendix C Extended Experiment Results
C.1 Additional Results on Token Quality Analysis and Frequency Learning
Figure 6: (a) Frequency and temporal token encoder ablation on TUAB. (b) & (c) presents Analysis of token quality across three TFM-Tokenizer variants and the neural tokenizer on IIIC. (b) Comparison of class-token uniqueness scores across all classes and (c) Class-wise token consistency analysis
Table 6: Token Utilization and class-token uniqueness comparison